# Accelerating Targeted Hard-Label Adversarial Attacks in Low-Query Black-Box Settings

### v1

@@ 4.2 Results
Figure 7: Comparison of median percentage decrease in $\ell^{2}$ distance across different architectures, with higher values indicating a more effective reduction method.
5 Conclusion
@@
Limitations. (i) Randomness: Similar to state-of-the-art methods, TEA involves inherent randomness. In the patch-based refinement step, positions and sizes of patches are chosen randomly, leading to varying results across runs. Nevertheless, the overall performance remains stable when averaged over the thousand image pairs tested. (ii) Low-query focus: TEA is optimized specifically for scenarios with limited query budgets. Beyond this low-query regime, the method’s progress plateaus in a narrower decision space. However, the adversarial examples generated at this stage serve as strong initializations for subsequent refinement by other existing techniques.
@@
We evaluate TEA exclusively on the publicly available ImageNet-1K validation set, which is distributed for non-commercial research and educational use under the ImageNet access agreement. The dataset contains no personally identifiable information, and we used it in accordance with the applicable license terms. As TEA reduces the number of required queries for targeted hard-label attacks in a low-query setting, it could be leveraged to more rapidly craft adversarial inputs against commercial or safety-critical vision systems. We encourage practitioners to adopt defense mechanisms that go beyond detecting frequency-based perturbations, explicitly incorporating checks for edge-informed distortions, and defenses attuned to other structural features, to guard against similar attacks.
Acknowledgements

### v3

@@ IV-B Results
Figure 8: Comparison of median percentage decrease in $\ell_{2}$ distance across different architectures — higher values indicate a more effective reduction method.
Source-Target Similarity Analysis. To examine how the performance of TEA depends on the similarity between source and target images, we group the $1000$ source-image pairs by structural similarity. For each pair $(x_{s},x_{t})$ we compute the SSIM score and sort all pairs by SSIM, partitioning them into ten equally sized bins. Within each bin, and model, we measure the percentage $\ell_{2}$ reduction at a fixed budget of $400$ queries. We then average this quantity over all pairs in the same bin and plot the resulting curves (Figure 9). We see that all methods have relatively identical performance variability, with improved performance across structurally similar source-target image pairs. This indicates that structural similarity does not favor TEA over the other methods.
ResNet50
ResNet101
VGG16
ViT
Figure 9: Effect of source–target structural similarity on attack performance. We group image pairs into ten bins by SSIM and report the average percentage decrease in $\ell_{2}$ distance after 400 queries. All methods benefit similarly from higher structural similarity.
Edge–Density Analysis. To assess how TEA behaves under different levels of structural detail in the source and target images, we perform an edge–density based stratification of the source–target pairs. For each image, we convert it to grayscale, compute Sobel gradients, and form the normalized gradient magnitude $\tilde{g}(p)\in[0,1]$ for each pixel $p$. Pixels with $\tilde{g}(p)>0.2$ are treated as edge pixels, and the edge density $d(x)$ is defined as the fraction of pixels classified as edges. We collect $d(x_{s})$ and $d(x_{t})$ for all source-target image pairs and set a global dense/sparse threshold $\tau_{\text{dense}}$ to be the median density over all $2000$ images, yielding $\tau_{\text{dense}}\approx 0.110242$ in our case. Images with $d(x)>\tau_{\text{dense}}$ are labeled dense, and the rest sparse. Each pair $(x_{s},x_{t})$ is thus assigned to one of four edge–pattern regimes: dense–sparse, dense–dense, sparse–sparse, or sparse–dense. For each regime, and model, we report the average percentage $\ell_{2}$ reduction (in %) at a fixed budget of $400$ queries. The resulting curves in Figure 10 show that performance variability is similar across methods, with higher success when moving towards a sparse source image than a dense one, and that TEA maintains a consistent advantage across all regimes.
ResNet50
ResNet101
VGG16
ViT
Figure 10: Effect of source–target edge density on attack performance. Each image is classified as dense or sparse based on a Sobel edge–density threshold, inducing four edge–pattern regimes (dense–dense, dense–sparse, sparse–dense, sparse–sparse). For each regime, and model, we report the average percentage decrease in $\ell_{2}$ distance after 400 queries, illustrating how performance varies with the edge richness of the source–target pairs.
Zero-Shot CLIP. To assess whether TEA also transfers to more modern vision-language models, we additionally evaluate it against a zero-shot CLIP classifier on the ImageNet validation set. We use the ViT-B/32 variant of CLIP [29, 13] in its standard zero-shot configuration. Following the usual protocol, we construct a linear zero-shot head by encoding multiple natural-language templates for each ImageNet class (e.g., “a photo of a {}.”, “a close-up photo of a {}.”) with the CLIP text encoder, normalizing and averaging the resulting embeddings to obtain a single prototype per class. Stacking these prototypes yields a fixed weight matrix, and logits are obtained by taking scaled dot products between normalized image features and these class prototypes. Our implementation relies on the official CLIP PyTorch code-base and its recommended preprocessing.
In Figure 11, we show the average percentage decrease in $\ell_{2}$ distance and the average $\ell_{2}$ distance as a function of the query budget. TEA retains a clear advantage, consistently achieving larger perturbation reductions than the baselines within the same limited query budget, indicating that its strategy remains effective even for modern zero-shot vision-language models.
Refer to caption Refer to caption
Figure 11: Comparison on a zero-shot CLIP (ViT-B/32) classifier on ImageNet. Left: average percentage decrease in $\ell_{2}$ distance against number of queries, with higher values indicating a more effective reduction method. Right: average $\ell_{2}$ distance against number of queries, with lower values indicating a more effective reduction method.
Other Datasets. To further assess the robustness and generality of TEA beyond ImageNet, we additionally evaluate all tested methods on two standard classification benchmark datasets: CIFAR-100 [23] and the Intel Image Classification dataset [22]. CIFAR-100 contains 100 object categories with lower-resolution images while the Intel dataset consists of six scene categories (e.g., buildings, forest, sea). In both cases, we again attack the four models, ResNet-50, ResNet-101, VGG-16, and ViT, and measure performance in terms of the relative $\ell_{2}$-distance reduction between the source and target images. Figures 12 and 13 report the average $\ell_{2}$-distance reduction at a fixed low-query budget of $400$ across all four models. Consistent with our results on the ImageNet dataset, TEA achieves larger perturbation reductions on both datasets, indicating that its edge-aware strategy remains effective across diverse data distributions and resolutions.
ResNet50
ResNet101
VGG16
ViT
Figure 12: Average $\ell_{2}$ distance reduction across different architectures in a low-query regime on the CIFAR-100 dataset.
ResNet50
ResNet101
VGG16
ViT
Figure 13: Average $\ell_{2}$ distance reduction across different architectures in a low-query regime on the Intel Image Classification dataset.
V Conclusion
@@
Limitations. TEA is specifically designed for the targeted, hard-label, low-query black-box setting, and its main advantage lies in rapidly improving performance under tight query budgets. When TEA is used merely as an initialization and followed by stronger high-query attacks in regimes where many queries are available, we observe that the final performance is comparable to that of the underlying state-of-the-art methods, rather than strictly better. In addition, while TEA preserves global edge structure and achieves low $\ell_{2}$ distances, the patch-wise updates can still introduce localized artifacts that are visible to human observers. This limitation is shared with other norm-bounded attacks in general, but in TEA the patch-based nature of the perturbations can make these local changes particularly noticeable.
@@
We evaluate TEA exclusively on the publicly available ImageNet-1K validation set, which is distributed for non-commercial research and educational use under the ImageNet access agreement. The dataset contains no personally identifiable information, and was used in accordance with the license terms. As TEA reduces the number of required queries for targeted hard-label attacks in a low-query setting, it could be leveraged to more rapidly craft adversarial inputs against commercial or safety-critical vision systems. We encourage practitioners to adopt defense mechanisms that go beyond detecting frequency-based perturbations, explicitly incorporating checks for edge-informed distortions, and defenses attuned to other structural features, to guard against similar attacks.
LLM Usage Considerations
LLMs were used for editorial purposes in this manuscript, and all outputs were inspected by the authors to ensure accuracy and originality. All technical ideas, experiments, analyses, and citations were conceived, implemented, and validated by the authors.
Acknowledgments